solve:
7/x - 2 = 4/x + 3
: solve:
7/x - 2 = 4/x + 3

Answer by pcsp_kumat(6)

7/x-2=4/x + 3 [transposing]
7/x -4/x= 3+2
3/x = 5 [cross multiplication]
3 = 5x
x = 3/5